# The largest 5th pivot may be the root of a 61st degree polynomial

### v1

@@ Abstract.
In 1988, Day and Peterson performed numerical optimization on $n=5$ with NPSOL and obtained a largest seen value of $4.1325$.... This same best value was reproduced by Gould with LANCELOT in 1991. We ran extensive comparable experiments with the modern software tool JuMP and also saw the same value $4.1325$.... While the combinatorial explosion of possibilities prevents us from knowing there may not be a larger maximum, we succeed in obtaining the exact mathematical value: the number $4.1325$... is exactly the root of a 61st degree polynomial provided in this work, and is a maximum given the equality constraints seen by JuMP. In light of the numerics, we pose the conjecture that this lower bound is indeed the maximum. We also apply this technique to $n=6$, 77, and $8$.
@@ 1. The maximum growth factor sequence for $n=1,\ldots,5$
This paper shows the exact value of $4.1325\ldots$ to have an unexpected form. We show that this number is the root of a 6161st degree polynomial with integer coefficients. We further conjecture that this lower bound is indeed the maximum (Conjecture 2.1). With the benefit of hindsight, we can see that this surprising form of an answer, the root of a 61st degree polynomial, could escape prediction. While the super-exponentially many possible configurations of tight constraints prevents us from fully proving this conjecture, we prove that this value is indeed the maximum when restricted to the set of constraints suggested by JuMP (Theorem 2.2). Our technique, involving non-linear optimization software, Grobner bases, and discriminants, provides a new approach to the growth factor problem. This approach can be readily applied to larger dimensions, and produces lower bounds for $n=6$, 77, and $8$, which we suppose are maximums as well. They are $5$ for $n=6$, a root of a sixth degree polynomial for $n=7$, and $8$ for $n=8$ (see Subsection 2.5). In addition, using a combination of mathematical analysis and interval arithmetic, we produce an improved upper bound of $4.84$ for the maximum growth factor when $n=5$ (Theorem 3.5), the first improvement upon Tornheim’s 1968 bound of $4\tfrac{17}{18}$ [20]. [^1] The recent mathematical improvement [1] on Wilkinson’s upper bound for the growth factor under complete pivoting does not improve upon Tornheim’s bound for $n=5$.
@@ 1.1. Gaussian elimination with complete pivoting
Here we briefly remind the reader of the definition of Gaussian elimination with complete pivoting and the associated growth factor. For a historical account of the growth factor problem in Gaussian elimination, we refer the reader to Higham’s classic text Accuracy and Stability of Numerical Algorithms [14, Chapter 9], [5, Section 1.1], and [21].
@@
Let $\|\cdot\|_{\max}$ be the largest magnitude entry of a matrix. The quantity
@@ 2.1. Our structure for $n=5$
We investigate in depth a specific elimination sequence structure that we have seen many times in our JuMP experiments, and also lead to the growth $4.1325\dots$ reported in experiments by [4, 7]. At the end of this investigation we will be able to conclude that the reported value of $4.1325\dots$ is exactly the root of a $61^{st}$ degree polynomial.
@@
Finally there is the introduction of the growth factor $g$ gives the equality $g={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bf 2\ \clubsuit}$. This may be rewritten as yet another polynomial expression ( Out[9]):
@@ 2.3. The role of the discriminant polynomial
Case 5: $(2x^{2}z+xz^{2}+xz-2x-2)$: Adding this to the Groebner Basis only allows $y$ to be -1 or $z$ to .     be 0 or $1$, all already ruled out. ∎
@@ 2.5. What about $n=6$, 77, and $8$?
When $n=7$, the JuMP computed matrix is more complex, but also has at least $n^{2}-1$ equality constraints, removing the need for any optimization, after the computation of a Groebner basis. Finally, we note that when $n=8$, numerical optimization suggests growth equal to $8$, a value known since the 1960’s, and achieved by Sylvester’s $8\times 8$ Hadamard matrix. However, interestingly, in 2024, we found non-Hadamard $8\times 8$ matrices that also achieve growth $8$ [5, Table 1].
@@ 3.3. Construction of covers
We note that in principle one can refine the side lengths of the boxes $B$ to be $2^{-l}$ for $l>4$ to obtain better approximations of $\mathcal{C}_{2.2}$ and $\mathcal{C}_{2.15}$, however this is expensive computationally due to the high dimensionality of the problem. The numbers $1882$ and $19511$ are not special, only that they are significantly smaller than the total number of such boxes $2^{36}$ and small enough for computation to be feasible.
@@
Fix a box $B$ in $S_{2.2}$ or $S_{2.15}$. We define a routine calc_range which given positive scalars $p_{2}$, $p_{3}$, a box $B$ and interval ranges for $u,v,y,x$, computes an interval containing $\min_{C\in B}\left\|p_{3}C-uv^{T}-p_{2}xy^{T}\right\|_{\max}$. The routine then true if the lower bound of the interval is strictly greater than $1$.

### v2

@@ Abstract.
In 1988, Day and Peterson performed numerical optimization on $n=5$ with NPSOL and obtained a largest seen value of $4.1325$.... This same best value was reproduced by Gould with LANCELOT in 1991. We ran extensive comparable experiments with the modern software tool JuMP and also saw the same value $4.1325$.... While the combinatorial explosion of possibilities prevents us from knowing whether there may not be a larger maximum, we succeed in obtaining the exact mathematical value: the number $4.1325$... is exactly the root of a 61st degree polynomial provided in this work, and is a maximum given the equality constraints seen by JuMP. In light of the numerics, we pose the conjecture that this lower bound is indeed the maximum. We also apply this technique to $n=6$, 77, and $8$.
@@ 1. The maximum growth factor sequence for $n=1,\ldots,5$
This paper shows the exact value of $4.1325\ldots$ to have an unexpected form. We show that this number is the root of a 6161st degree polynomial with integer coefficients. We further conjecture that this lower bound is indeed the maximum (Conjecture 2.1). With the benefit of hindsight, we can see that this surprising form of an answer, the root of a 61st degree polynomial, could escape prediction. While the super-exponentially many possible configurations of tight constraints prevents us from fully proving this conjecture, we prove that this value is indeed tight when restricted to the set of constraints suggested by JuMP (Theorem 2.2). Our technique, involving non-linear optimization software, Grobner bases, and discriminants, provides a new approach to the growth factor problem. This approach can be readily applied to larger dimensions, and produces lower bounds for $n=6$, 77, and $8$, which we suppose are maximums as well. They are $5$ for $n=6$, a root of a sixth degree polynomial for $n=7$, and $8$ for $n=8$ (see Subsection 2.5). In addition, using a combination of mathematical analysis and interval arithmetic, we produce an improved upper bound of $4.84$ for the maximum growth factor when $n=5$ (Theorem 3.5), the first improvement upon Tornheim’s 1968 bound of $4\tfrac{17}{18}$ [20]. [^1] The recent mathematical improvement [1] on Wilkinson’s upper bound for the growth factor under complete pivoting does not improve upon Tornheim’s bound for $n=5$.
@@ 1.1. Gaussian elimination with complete pivoting
Here we briefly remind the reader of the definition of Gaussian elimination with complete pivoting and the associated growth factor. For a historical account of the growth factor problem in Gaussian elimination, we refer the reader to Higham’s classic text Accuracy and Stability of Numerical Algorithms [14, Chapter 9], [5, Section 1.1], and [21]. Here and in what follows, we restrict ourselves to real $n\times n$ matrices.
@@
Let $\|\cdot\|_{\max}$ denote the largest magnitude entry of a matrix. The quantity
@@ 2.1. Our structure for $n=5$
We investigate in depth a specific elimination sequence structure that we have seen many times in our JuMP experiments and that leads to the growth $4.1325\dots$ reported in experiments by [4, 7]. At the end of this investigation we will be able to conclude that the reported value of $4.1325\dots$ is exactly the root of a $61^{st}$ degree polynomial.
@@
Finally there is the introduction of the growth factor $g$ that gives the equality $g={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bf 2\ \clubsuit}$. This may be rewritten as yet another polynomial expression ( Out[9]):
@@ 2.3. The role of the discriminant polynomial
Case 5: $(2x^{2}z+xz^{2}+xz-2x-2)$: Adding this to the Groebner Basis only allows $y$ to be 1-1 or $z$ to .     be 0 or $1$, all already ruled out. ∎
@@ 2.5. What about $n=6$, 77, and $8$?
When $n=7$, the JuMP computed matrix is more complicated, but also has at least $n^{2}-1$ equality constraints, removing the need for any optimization, after the computation of a Groebner basis. Finally, we note that when $n=8$, numerical optimization suggests growth equal to $8$, a value known since the 1960’s, and achieved by Sylvester’s $8\times 8$ Hadamard matrix. However, interestingly, in 2024, we found non-Hadamard $8\times 8$ matrices that also achieve growth $8$ [5, Table 1].
@@ 3.3. Construction of covers
We note that in principle one can refine the side lengths of the boxes $B$ to be $2^{-l}$ for $l>4$ to obtain better approximations of $\mathcal{C}_{2.2}$ and $\mathcal{C}_{2.15}$. However this is expensive computationally due to the high dimensionality of the problem. The numbers $1882$ and $19511$ are not special, only that they are significantly smaller than the total number of such boxes $2^{36}$ and small enough for computation to be feasible.
@@
Fix a box $B$ in $S_{2.2}$ or $S_{2.15}$. We define a routine calc_range which given positive scalars $p_{2}$, $p_{3}$, a box $B$ and interval ranges for $u,v,y,x$, computes an interval containing $\min_{C\in B}\left\|p_{3}C-uv^{T}-p_{2}xy^{T}\right\|_{\max}$. The routine then returns true if the lower bound of the interval is strictly greater than $1$.